A student is taking a true/false exam with 16 questions. If he guesses on each question, what is the probability he will get at least 14 questions correct?
Found 2 solutions by stanbon, nyc_function:
Answer by stanbon(75887)

A student is taking a true/false exam with 16 questions. If he guesses on each question, what is the probability he will get at least 14 questions correct?
-------------------------------------------
P(14<= x <=16)
= 1 - 16C14(1/2)^14*(1/2)^2
---
= 1 - 120(1/2)^16
---
= 1 - 0.001831
---
= 0.9982
=============
Cheers,
Stan H.

Answer by nyc_function(2741)

There are 16 questions. For each question, the student has two answers to select from.
TIP: 16 x 2 = 32 = total choices
Let p = probability
p(14 correct) = 14/total choices
Can you do it now?